OpenStudy (anonymous):

Contest participants are each assigned a ticket in one of the following colors: red, blue, or orange. There are 8 red tickets assigned, 20 blue tickets assigned, and 12 orange tickets assigned. If one participant is to be selected at random, what is the likelihood that the selected person's assigned ticket will be orange?

OpenStudy (kropot72):

The probability that the selected person's ticket will be orange is given by: \[P(orange)=\frac{quantity\ of\ orange\ tickets}{total\ quantity\ of\ all\ 3\ colors}=you\ can\ calculate\]

OpenStudy (anonymous):

I get 0.3 as the final answer and I know that the correct answer is 3/10 but how 3/10?

OpenStudy (kropot72):

Quantity of orange tickets is 12. Total quantity of all 3 colors = 8 + 20 + 12 = 40  O.3 is the decimal equivalent of 3/10